# Swarm-based gradient descent meets simulated annealing

### v1

@@ Abstract.
We introduce a novel method for non-convex optimization which is at the interface between the swarm-based gradient-descent (SBGD) [21, 28], and Simulated Annealing (SA) [5, 18, 11]. We follow the methodology of SBGD in which a swarm of agents, each identified with a position, $\mathbf{x}$ and mass $m$, explores the ambient space. The agents proceed in gradient descent direction, and are subject to Brownian motion with annealing-rate dictated by a decreasing function of their mass. Thus, instead of the SA protocol for time-decreasing temperature, we let the swarm decide how to ‘cool down’ agents, depending on their accumulated mass over time. The dynamics of masses is coupled with the dynamics of positions: agents at higher ground transfer (part of) their mass to those at lower ground. Consequently, the swarm is dynamically divided between heavier, cooler agents viewed as ‘leaders’ and lighter, warmer agents viewed as ‘explorers’. Mean-field convergence analysis and benchmark optimizations demonstrate the effectiveness of the swarm-based method as a multi-dimensional global optimizer.
@@
Acknowledgments. Research was supported by NSF grant DMS-2012292 (MG), by NSF grant DMS-2308440 and ONR grant N00014-21-1-2140 (QL) and by ONR grant N00014-2112773 (ET). ET thanks the Fondations Sciences Mathématiques des Paris (FSMP) and LJLL at the Sorbonne University for the support and hospitality.
@@ 1. Introduction
where $\overline{F}^{N}_{t}$ is the mass-weighted average of the $N$ agents, $\overline{F}^{N}_{t}:=\frac{\sum_{j=1}^{N}m^{j}_{t}F(\mathbf{x}^{j}_{t})}{\sum_{j=1}^{N}m^{j}_{t}}$. We tune the annealing rates as a decreasing function of the mass, $\sigma^{j}_{t}=\sigma(m^{j}_{t})$. Thus, our protocol for annealing-rate prevents lighter agents from being trapped in basins of local minima while ‘cools down’ heavier agents at the basin of attraction of global minimum of $F$. This is where the swarm-based gradient descent [21] meets Simulated Annealing (SA), [18, 10, 5], except that the protocol for “cooling process” is different: instead of an explicit recipe for decreasing the temperature as done in SA, our swarm-based approach let different agents at a lower ground cool down, depending on their increasing mass $\sigma^{j}_{t}=\sigma(m^{j}_{t})$.
Why provisional minimum? We refer to the mass-weighted average $\overline{F}^{N}_{t}:=\frac{\sum_{j=1}^{N}m^{j}_{t}F(\mathbf{x}^{j}_{t})}{\sum_{j=1}^{N}m^{j}_{t}}$ as the provisional minimum. To clarify why provisional minimum is needed, we note that according to the mass equation above, agents that are above the provisional minimum will shed a fraction of their mass — masses that are transferred to agents that are below the provisional minimum. Since the total mass remains constant, $\sum_{j}\textnormal{d}m^{j}_{t}/\textnormal{d}t=0$, masses tend to concentrate with agents near or below the provisional minimum, and it therefore makes sense to adjust the annealing rate, $\sigma^{j}_{t}$, as a decreasing function of the mass,
@@
Given this adjustment of mass-dependent annealing, it is expected that for large crowd of agents, $N\gg 1$, the provisional minimum will approach the global minimum,
@@
Indeed, our main result states that, under the appropriate assumptions outlined in Theorem 2.3 below, this is true.
The method described above lies at the interface of the deterministic swarm-based gradient descent (SBGD) approach and the stochastic-based SA. We recall that the SBGD is governed by a deterministic system which governs the swarm of $N$-agents
(1.1) $\left\{\begin{split}\frac{\textnormal{d}}{\textnormal{d}t}\mathbf{x}^{j}(t)&=-\alpha(m^{j}(t))\nabla F(\mathbf{x}^{j}(t)),\\ \frac{\textnormal{d}}{\textnormal{d}t}m^{j}(t)&=-\big{(}F(\mathbf{x}^{j}(t))-\overline{F}^{N}(t)\big{)}m^{j}(t),\end{split}\right.\qquad j=1,2,\ldots,N.$
The key feature advocated in (1.1), is swarm dynamics of both — positions and masses embedded in ${\mathbb{R}}^{d}\times{\mathbb{R}}_{+}$. The additional dimension of mass serves as a platform for communication among agents, encoded in the provisional minimum, which in [21] was taken as the minimum of the crowd at the given time, $\overline{F}^{N}(t)=\min_{j}F(\mathbf{x}^{j}(t))$. The same swarm-based methodology is used in this work, except two distinct features. First, the provisional minimum is given here by the weighted average $\overline{F}^{N}_{t}:=\frac{\sum_{j}m^{j}_{t}F(\mathbf{x}^{j}_{t})}{\sum_{j}m^{j}_{t}}$ which eventually is expected to approach the global minimum, $\min_{j}F(\mathbf{x}^{j}(t))$, used in [21]. The second and more essential distinction, is how masses are being used to adjust the dynamics of positions along the gradient descent: in SBGD, it is the time step which is adjusted as a decreasing function of the mass, $\alpha=\alpha(m^{j}(t))$. In the swarm-based optimization proposed in this work, masses are used to adjust the annealing rate, $\sigma=\sigma(m^{j}_{t})$, as decreasing function of mass,
(1.2) $\left\{\begin{split}\textnormal{d}\mathbf{x}^{j}_{t}&=-\nabla F(\mathbf{x}^{j}_{t})\textnormal{d}t+\sqrt{2\sigma(m^{j}_{t})}\textnormal{d}W^{j}_{t},\\ \textnormal{d}m^{j}_{t}&=-\big{(}F(\mathbf{x}^{j}_{t})-\overline{F}^{N}_{t}\big{)}m^{j}_{t}\textnormal{d}t,\quad\overline{F}^{N}_{t}:=\frac{\sum_{j=1}^{N}m^{j}_{t}F(\mathbf{x}^{j}_{t})}{\sum_{j=1}^{N}m^{j}_{t}},\end{split}\right.\qquad j=1,2,\ldots,N.$
@@
The SBGD method (1.1) employs the actual minimum, $\overline{F}(t)=\min_{j}F(\mathbf{x}^{j}(t))$ as its provisional minimum [21, 28]. This, however, is not amenable to mean-field limit that we will pursue in Section 2.1. The mean-field limit requires us to tour the whole landscape, ending with
@@
which is already the global minimum we are looking for. Instead, the use of mass-weighted average $\overline{F}^{N}_{t}$ in (1.2) as a provisional minimum, admits a mean-field interpretation which is shown to be driven towards the same desired global minimum.
@@ 1.1. In-swarm communication combined with stochastic search
Application of deterministic swarm-based methods reveals that such methods succeed in the accuracy sense but fail in the probability sense, namely — when these methods succeed, then they find the global optimizer, yet there is non-zero probability for the methods to fail to do so. In contrast, stochastic methods succeed in the probability sense but fail in the accuracy sense, namely — such methods can always find the global basin, but they do not necessarily return a true global optimizer. In this context, our swarm-based optimization with adjusted annealing rate  takes advantage of randomness in (1.2)1 combined with communication-based swarming in (1.2)2, with the aim of succeeding in both the probability and the accuracy sense. This is achieved by a dynamic process that combines swarm communication and randomness, as described below.
The essential role of communication in SBGD was already emphasized in [21, 28]. This is amplified in the present context, upon setting $\sigma(\cdot)\equiv 0$ in (1.2)1. One ends up with a crowd of $N$ independent agents driven by gradient descent,
@@
The success of such a non-communicating crowd in exploring the ambient landscape is significantly worse than that of a communication-based crowd. In Section  4, we show that the method fails in the probability sense.
Communication is encoded in the provisional minimum, where agents ‘communicate’ their height relative to $\overline{F}^{N}_{t}$. In our swarm-based optimization with adjusted annealing rate  (1.2), $\overline{F}^{N}_{t}$ is taken as the weighted average height, weighted by the different masses. Consequently, there is a dynamic distinction between agents above the provisional minimum and agents below or at the level of the provisional minimum. Agents above the provisional minimum carry lighter mass and therefore explore the ambient landscape with a relatively large Brownian motion; these are viewed as the explorers of the crowd. The heavier agents — those that are below or at the level of the provisional minimum, evolve in the gradient direction with a smaller Brownian motion and are viewed as leaders of the crowd. Of course, once an exploring light agent ‘hits’ a new lower ground at, say, $\mathbf{x}^{j}_{t}$ where $F(\mathbf{x}^{j}_{t})\ll-\overline{F}^{N}_{t}$, it is expected to accumulate more mass from the higher agents, to cool down, and eventually to become a leader, driving the crowd to a new lower point.
We note that by fixing a constant annealing rate, (1.2)1 with $\sigma(m^{j}_{t})\equiv\sigma>0$, we also encounter a problem. In this case, we recover Langevin Monte Carlo
@@
whose long time invariant measure suggests $\mathbf{x}\propto e^{-F(\mathbf{x})/\sigma}$. It admits the proper convergence in the zero-variance limit, $\mathbf{x}\stackrel{{\scriptstyle\sigma\rightarrow 0}}{{\longrightarrow}}\operatorname*{\textnormal{argmin}}_{\mathbf{x}}F(\mathbf{x})$, but, as we show in Section 4 below, the resulting method fails in the accuracy sense for any finite $\sigma$. Thus, the dynamic process of relabeling explorers and leaders, coupled through a mass-dependent annealing rate, $\sigma^{j}_{t}=\sigma(m^{j}_{t})$, is at the heart of matter.
In Section 5 we prove global convergence of (1.2) in both — sense of probability and sense of accuracy. This is achieved in a two-step argument: first, large-crowd, mean-field analysis summarized in Theorem 2.1 below, which allows us to pass from finitely many Stochastic-ODEs governing the particle system (1.2), to the limiting dynamics for the probability measure, $\mu_{t}(\mathbf{x},m)$, governed by the limiting PDE (2.4); and second, the large-time behavior of the latter, summarized in Theorem 2.2, implies that in this mean-field limit, the PDE in long time returns the the provisional minimum as the global minimum. Combining these results we conclude in Theorem 2.3 the claimed convergence $\mathop{\liminf}_{t\rightarrow\infty}\lim_{N\rightarrow\infty}\overline{F}^{N}_{t}=F_{*}$ with a precise convergence rate of polynomial type. Further, in Section 3 we derive the corresponding macroscopic description of the system.
@@ 1.2. The swarm-based optimization with adjusted annealing rate
The exchange of mass encoded in (1.2) assigns smaller masses to agents with higher values, while the decreasing property of $\sigma(m)$ ensures that these agents experience larger random perturbations. As a result, lighter agents serve as explorers, while heavier agents serve as leaders. As noted in [21], upon normalization of the total mass, $\sum_{j}m^{j}_{t}\equiv\sum_{j}m^{j}_{0}=1$, masses can be interpreted as ‘probabilities of finding a global minimum’; heavier agents have larger probability in doing so. The computation proceeds by discretization of (1.2) using Euler-Maruyama’s formula with a proper time step $h>0$,
@@
for $j=1,2,...,N$ and $\boldsymbol{\xi}^{j}\sim\mathcal{N}(0,\mathbf{I}_{d})$ i.i.d for all $j$. The detailed algorithm is summarized in Algorithm 1.
The performance of the algorithm is presented in Section 6. This follows the analytical study in Sections 2 and 5 in which we state, and respectively, prove the main analytical results of the paper, regarding the large-crowd, large-time dynamics associated with (1.2).
Algorithm 1 Swarm-based stochastic gradient descent
@@ 1.3. Related work
It is here that we observe the main difference between our swarm-based optimization with adjusted annealing rate  (1.2) and the CBO (1.4): while the latter requires a user tuning of parameter $\lambda$ in $\overline{F}^{\lambda}$, the swarm-based framework of [21] provides us with an adaptive mechanism which dynamically adjusts the provisional minimum based on the mass distribution in different parts of the landscape, $\overline{F}^{N}_{t}$, and the corresponding mean-field ¯$\overline{\mathscr{F}}^{\mu}_{t}$, outlined in (2.4) below.
@@ 2.1. From empirical distribution to mean-field
We define the empirical distribution $\mu^{N}_{t}$, which records the ensemble distribution of $\{\mathbf{x}^{j}_{t},m^{j}_{t}\}_{j=1}^{N}$ that obey (1.2), as a measure on the feature space $(\mathbf{x},m)$,
@@
Differentiating, calling Itô’s formula, and using (1.2) yield
@@
Observe that ¯$\overline{\mathscr{F}}^{\mu}_{t}$ denotes the provisional minimum associated with mean-field $\mu$, which is to be distinguished from the discrete provisional minimum in (2.2), denoted $\overline{F}^{N}_{t}$. We note that, similar to the discrete mass conservation in (1.2), $\frac{1}{N}\sum_{j}m^{j}_{t}\equiv 1$, the total mass in (2.4a) is conserved and equals one
@@
Assume that Assumption 2.1 holds. Let $\mu_{t}=\mu_{t}(\mathbf{x},m)$ be the mean-field solution of (2.4a) and let $\mu^{N}_{t}=\frac{1}{N}\sum_{j=1}^{N}\delta_{\mathbf{x}^{j}_{t}}(\mathbf{x})\otimes\delta_{m^{j}_{t}}(m)$ be the empirical distribution associated with the ensemble of swarm-based solutions  (1.2) subject to compatible initial data so that $\{\mathbf{x}^{j}_{0}\,,m^{j}_{0}\}$ are i.i.d. samples of $\mu_{t=0}(\mathbf{x},m)$. Then $\mu_{t}$ and $\mu^{N}_{t}$ are close in the Wasserstein sense:
(2.5) $\lim_{N\rightarrow\infty}\mathbb{E}\left[W_{2}\left(\mu_{t},\mu^{N}_{t}\right)\right]=0\,,$

### v2

@@ Abstract.
We introduce a novel method, called Swarm-based Simulated Annealing (SSA), for non-convex optimization which is at the interface between the swarm-based gradient-descent (SBGD) [21, 29], and Simulated Annealing (SA) [5, 18, 11]. Similar to SBGD, we introduce a swarm of agents, each identified with a position, $\mathbf{x}$ and mass $m$, to explore the ambient space. Similar to SA, the agents proceed in the gradient descent direction, and are subject to Brownian motion. The annealing rate, however, is dictated by a decreasing function of their mass. As a consequence, instead of the SA protocol for time-decreasing temperature, here the swarm decides how to ‘cool down’ agents, depending on their own accumulated mass. The dynamics of masses is coupled with the dynamics of positions: agents at higher ground transfer (part of) their mass to those at lower ground. Consequently, the resulting SSA optimizer is dynamically divided between heavier, cooler agents viewed as ‘leaders’ and lighter, warmer agents viewed as ‘explorers’. Mean-field convergence analysis and benchmark optimizations demonstrate the effectiveness of the SSA method as a multi-dimensional global optimizer.
@@
Acknowledgments. Research was supported by NSF grant DMS-2012292 (MG), by NSF grant DMS-2308440 and ONR grant N00014-21-1-2140 (QL) and by ONR grants N00014-2112773, N00014-2412659 (ET). ET thanks the Fondations Sciences Mathématiques des Paris (FSMP) and LJLL at the Sorbonne University for the support and hospitality.
@@ 1. Introduction
where $\overline{F}^{N}_{t}$ is the mass-weighted average of the $N$ agents, $\overline{F}^{N}_{t}:=\frac{\sum_{j=1}^{N}m^{j}_{t}F(\mathbf{x}^{j}_{t})}{\sum_{j=1}^{N}m^{j}_{t}}$. We tune the annealing rates as a decreasing function of the mass, $\sigma^{j}_{t}=\sigma(m^{j}_{t})$. Thus, our protocol for annealing-rate prevents lighter agents from being trapped in basins of local minima while ‘cools down’ heavier agents at the basin of attraction of global minimum of $F$. This is where the Swarm-Based Gradient Descent (SBGD) [21] meets Simulated Annealing (SA), [18, 10, 5], except that the protocol for “cooling process” is different: instead of an explicit recipe for decreasing the temperature as is done in SA, our swarm-based approach lets different agents at a lower ground cool down, depending on their increasing mass $\sigma^{j}_{t}=\sigma(m^{j}_{t})$.
Why provisional minimum? We refer to the mass-weighted average
(1.1) $\overline{F}^{N}_{t}:=\frac{\sum_{j=1}^{N}m^{j}_{t}F(\mathbf{x}^{j}_{t})}{\sum_{j=1}^{N}m^{j}_{t}}$
as the provisional minimum. To clarify why provisional minimum is needed, we note that according to the mass equation above, agents that are above the provisional minimum will shed a fraction of their mass — masses that are transferred to agents that are below the provisional minimum. Since the total mass remains constant, $\sum_{j}\textnormal{d}m^{j}_{t}/\textnormal{d}t=0$, masses tend to concentrate with agents near or below the provisional minimum, and it therefore makes sense to adjust the annealing rate, $\sigma^{j}_{t}$, as a decreasing function of the mass,
@@
Given this adjustment of mass-dependent annealing, it is expected that for a large crowd of agents, $N\gg 1$, (1.1) will approach the global minimum,
@@
Indeed, our main result states that, under the appropriate assumptions (in particular, the assumption of uniqueness of global minimizer made in Assumption 2.2 below), this is true.
The method described above lies at the interface of the deterministic SBGD approach and the stochastic-based SA. We recall that the SBGD is governed by a deterministic system which governs the swarm of $N$-agents, and for $j=1,2,\ldots,N$
(SBGD) $\frac{\textnormal{d}}{\textnormal{d}t}\mathbf{x}^{j}(t)=-\alpha(m^{j}(t))\nabla F(\mathbf{x}^{j}(t)),\quad\frac{\textnormal{d}}{\textnormal{d}t}m^{j}(t)=-\big{(}F(\mathbf{x}^{j}(t))-\overline{F}^{N}(t)\big{)}m^{j}(t)\,.$
The key feature advocated in (SBGD) is swarm dynamics of both positions and masses embedded in ${\mathbb{R}}^{d}\times{\mathbb{R}}_{+}$. The additional dimension of mass serves as a platform for communication among agents, encoded in the provisional minimum, which in [21] was taken as the minimum of the crowd at the given time, $\overline{F}^{N}(t)=\min_{j}F(\mathbf{x}^{j}(t))$. The same swarm-based methodology is used in this work, except for two distinct features. First, the provisional minimum is given here by (1.1) which eventually is expected to approach the global minimum, $\min_{j}F(\mathbf{x}^{j}(t))$, used in [21]. The second and more essential distinction, is how masses are being used to adjust the dynamics of positions along the gradient descent: in SBGD, it is the time step which is adjusted as a decreasing function of the mass, $\alpha=\alpha(m^{j}(t))$. In our new proposed algorithm, masses are used to adjust the annealing rate, $\sigma=\sigma(m^{j}_{t})$, as a decreasing function of mass. This new algorithm is named Swarm-based Simulated Annealing (SSA) optimizer:
(SSA) $\left\{\begin{split}\textnormal{d}\mathbf{x}^{j}_{t}&=-\nabla F(\mathbf{x}^{j}_{t})\textnormal{d}t+\sqrt{2\sigma(m^{j}_{t})}\textnormal{d}W^{j}_{t},\\ \textnormal{d}m^{j}_{t}&=-\big{(}F(\mathbf{x}^{j}_{t})-\overline{F}^{N}_{t}\big{)}m^{j}_{t}\textnormal{d}t,\end{split}\right.\qquad j=1,2,\ldots,N.$
@@
The SBGD method (SBGD) employs the actual minimum, $\overline{F}(t)=\min_{j}F(\mathbf{x}^{j}(t))$ as its provisional minimum [21, 29]. This, however, is not amenable to the mean-field limit analysis that we will pursue in Section 2.1. The mean-field limit requires us to tour the whole landscape, ending with
@@
which is already the global minimum we are looking for. Instead, the use of mass-weighted average $\overline{F}^{N}_{t}$ in (SSA) as a provisional minimum, admits a mean-field interpretation which is shown to be driven towards the same desired global minimum.
@@ 1.1. In-swarm communication combined with stochastic search
Application of deterministic swarm-based methods reveals that such methods succeed in the accuracy sense but fail in the probability sense, — namely when these methods succeed, then they find the global optimizer, yet there is non-zero probability for the methods to fail to do so. In contrast, stochastic methods succeed in the probability sense but fail in the accuracy sense, — namely such methods can always find the global basin, but they do not necessarily return a true global optimizer. In this context, our swarm-based optimization with adjusted annealing rate  takes advantage of randomness in (SSA)1 combined with communication-based swarming in (SSA)2, with the aim of succeeding in both the probability and the accuracy sense. This is achieved by a dynamic process that combines swarm communication and randomness, as described below.
The essential role of communication in SBGD was already emphasized in [21, 29]. This is amplified in the present context, upon setting $\sigma(\cdot)\equiv 0$ in (SSA)1. One ends up with a crowd of $N$ independent agents driven by gradient descent,
@@
The success of such a non-communicating crowd in exploring the ambient landscape is significantly worse than that of a communication-based crowd. In Section 4, we show that the method fails in the probability sense.
Communication is encoded in the provisional minimum, where agents ‘communicate’ their height relative to $\overline{F}^{N}_{t}$. In our swarm-based optimization with adjusted annealing rate  (SSA), $\overline{F}^{N}_{t}$ is taken as the weighted average height, weighted by the different masses. Consequently, there is a dynamic distinction between agents above the provisional minimum and agents below or at the level of the provisional minimum. Agents above the provisional minimum carry lighter mass and therefore explore the ambient landscape with a relatively large Brownian motion; these are viewed as the explorers of the crowd. The heavier agents — those that are below or at the level of the provisional minimum, evolve in the gradient direction with a smaller Brownian motion and are viewed as leaders of the crowd. Of course, once an exploring light agent ‘hits’ a new lower ground at, say, $\mathbf{x}^{j}_{t}$ where $F(\mathbf{x}^{j}_{t})\ll-\overline{F}^{N}_{t}$, it is expected to accumulate more mass from the higher agents, to cool down, and eventually to become a leader, driving the crowd to a new lower point.
We note that by fixing a constant annealing rate, (SSA)1 with $\sigma(m^{j}_{t})\equiv\sigma>0$, we also encounter a problem. In this case, we recover Langevin Monte Carlo
@@
whose long time invariant measure suggests $\mathbf{x}\propto e^{-F(\mathbf{x})/\sigma}$. It admits the proper convergence in the zero-variance limit, $\mathbf{x}\stackrel{{\scriptstyle\sigma\rightarrow 0}}{{\longrightarrow}}\operatorname*{\textnormal{argmin}}_{\mathbf{x}}F(\mathbf{x})$, but, as we show in Section 4 below, the resulting method fails in the accuracy sense for any finite $\sigma$. Thus, the dynamic process of relabeling explorers and leaders, coupled through a mass-dependent annealing rate, $\sigma^{j}_{t}=\sigma(m^{j}_{t})$, is at the heart of the matter.
In Section 5 we prove global convergence of (SSA) in both, the sense of probability and the sense of accuracy. This is achieved in a two-step argument: first, we conduct a large-crowd, mean-field analysis summarized in Theorem 2.1 below, which allows us to pass from finitely many Stochastic-ODEs governing the particle system (SSA), to the limiting dynamics for the probability measure, $\mu_{t}(\mathbf{x},m)$, governed by the limiting PDE (2.4); and second, the large-time behavior of the latter, summarized in Theorem 2.2, implies that in this mean-field limit, the PDE in long time returns the the provisional minimum as the global minimum. Combining these results we conclude in Theorem 2.3 the claimed convergence $\mathop{\liminf}_{t\rightarrow\infty}\lim_{N\rightarrow\infty}\overline{F}^{N}_{t}=F_{*}$ with a precise convergence rate of polynomial type. Further, in Section 3 we derive the corresponding macroscopic description of the system.
@@ 1.2. The swarm-based optimization with adjusted annealing rate
The exchange of mass encoded in (SSA) assigns smaller masses to agents with higher values, while the decreasing property of $\sigma(m)$ ensures that these agents experience larger random perturbations. As a result, lighter agents serve as explorers, while heavier agents serve as leaders. As noted in [21], upon normalization of the total mass, $\sum_{j}m^{j}_{t}\equiv\sum_{j}m^{j}_{0}=1$, masses can be interpreted as ‘probabilities of finding a global minimum’; heavier agents have larger probability in doing so. The computation proceeds by discretization of (SSA) using Euler-Maruyama’s formula with a proper time step $h>0$,
@@
for $j=1,2,...,N$ and $\boldsymbol{\xi}^{j}\sim\mathcal{N}(0,\mathbf{I}_{d})$ i.i.d for all $j$. The detailed algorithm is summarized in Algorithm 1. The mass is conserved throughout the algorithm. According to Line 6, the update gives: $\sum_{j=1}^{N}m^{j}_{n+1}=\sum_{j=1}^{N}m^{j}_{n}-h\sum_{j=1}^{N}m^{j}_{n}\left(F(\mathbf{x}^{j}_{n})-\overline{F}_{n}\right)=\sum_{j=1}^{N}m^{j}_{n}$, where the second summation term becomes zero owing to the definition of $\overline{F}_{n}$ shown in Line 3 in the algorithm.
The performance of the algorithm is presented in Section 6. This follows the analytical study in Sections 2 and 5 in which we state, and respectively, prove the main analytical results of the paper, regarding the large-crowd, large-time dynamics associated with (SSA).
Algorithm 1 Swarm-based Simulated Annealing (SSA)
@@ 1.3. Related work
It is here that we observe the main difference between our swarm-based optimization with adjusted annealing rate  (SSA) and the CBO (1.3): while the latter requires a user tuning of parameter $\lambda$ in $\overline{F}^{\lambda}$, the swarm-based framework of [21] provides us with an adaptive mechanism which dynamically adjusts the provisional minimum based on the mass distribution in different parts of the landscape, $\overline{F}^{N}_{t}$, and the corresponding mean-field ¯$\overline{\mathscr{F}}^{\mu}_{t}$, outlined in (2.4) below.
@@ 2.1. From empirical distribution to mean-field
We define the empirical distribution $\mu^{N}_{t}$, which records the ensemble distribution of $\{\mathbf{x}^{j}_{t},m^{j}_{t}\}_{j=1}^{N}$ that obey (SSA), as a measure on the feature space $(\mathbf{x},m)$,
@@
Differentiating, recalling Itô’s formula, and using (SSA) yield
@@
Observe that ¯$\overline{\mathscr{F}}^{\mu}_{t}$ denotes the provisional minimum associated with mean-field $\mu$, which is to be distinguished from the discrete provisional minimum in (2.2), denoted $\overline{F}^{N}_{t}$. We note that, similar to the discrete mass conservation in (SSA), $\frac{1}{N}\sum_{j}m^{j}_{t}\equiv 1$, the total mass in (2.4a) is conserved and equals one
@@
Assume that Assumption 2.1 holds. Let $\mu_{t}=\mu_{t}(\mathbf{x},m)$ be the mean-field solution of (2.4a) and let $\mu^{N}_{t}=\frac{1}{N}\sum_{j=1}^{N}\delta_{\mathbf{x}^{j}_{t}}(\mathbf{x})\otimes\delta_{m^{j}_{t}}(m)$ be the empirical distribution associated with the ensemble of swarm-based solutions  (SSA) subject to compatible initial data so that $\{\mathbf{x}^{j}_{0}\,,m^{j}_{0}\}$ are i.i.d. samples of $\mu_{t=0}(\mathbf{x},m)$. Then $\mu_{t}$ and $\mu^{N}_{t}$ are close in the Wasserstein sense:
(2.5) $W_{2}\left(\mu_{t},\mu^{N}_{t}\right)\rightarrow 0,\quad\text{in probability as $N\rightarrow\infty$}\,,$